# Decomposition-Based Optimal Bounds for Privacy Amplification via Shuffling

### v5

@@ 8 Limitation
However, when the adjacency relation is defined over the full product domain $\mathbb{X}^{1}\times\mathbb{X}^{2}\times\dots\times\mathbb{X}^{m}$, one must also consider neighboring pairs that differ in only a subset of coordinates. If all $\mathcal{R}_{i}$ are identical, we need to separately consider adjacency relations with Hamming distances $1,2,\dots,m$. For example, in the joint composition of three $\varepsilon_{0}/4$-DP $k$-RR mechanisms, we must evaluate the GPARV for the following input pairs:
$\big{(}(0,0,0,0),(1,1,1,1)\big{)},\quad\big{(}(0,0,0,0),(0,1,1,1)\big{)},\quad\big{(}(0,0,0,0),(0,0,1,1)\big{)},\quad\big{(}(0,0,0,0),(0,0,0,1)\big{)}.$ .
@@ A.1 Binary local hash
Let $\mathcal{X}=[D]$. The Binary Local Hashing (BLH) mechanism is defined as $\mathsf{BLH}(x)=(h,\mathsf{RR}(h(x)))$, where $h:[D]\to[2]$ is a uniformly random function drawn from =$\mathcal{H}=\{f:[D]\to[2]\}$, and $\mathsf{RR}$ denotes the 2-ary randomized response mechanism (see Example 4.1 for the definition) [WBLJ17].
@@
In summary, the blanket distribution of BLH is “nearly” uniform distribution over ×$\mathcal{H}\times[2]$:
@@
$S_{r_{0},r_{1}}=\left\{y\ \middle|\ \frac{\mathcal{R}(x_{1}^{0})(y)}{\inf_{x}\mathcal{R}(x)(y)}=r_{0},\frac{\mathcal{R}(x_{1}^{1})(y)}{\inf_{x}\mathcal{R}(x)(y)}=r_{1}\right\}.$
@@
$S_{e^{\varepsilon_{0}},1}$ =$=\{(h,y)\mid h(x_{1}^{0})=y\wedge h(x_{1}^{1})\neq y\},$
$S_{1,e^{\varepsilon_{0}}}$ =$=\{(h,y)\mid h(x_{1}^{0})\neq y\wedge h(x_{1}^{1})=y\},$
$S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}}$ =$=\{(h,y)\mid h(x_{1}^{0})=y\wedge h(x_{1}^{1})=y\}-\{(h,y)\mid h=h^{*}\vee h=\bar{h}^{*}\},$
$S_{1,1}$ =$=\{(h,y)\mid h(x_{1}^{0})\neq y\wedge h(x_{1}^{1})\neq y\}\cup\{(h,y)\mid h=h^{*}\vee h=\bar{h}^{*}\}.$
@@
$p$ $=\gamma\omega(S_{e^{\varepsilon_{0}},1})$ $\quad$=$\quad$= $\frac{1}{2(e^{\varepsilon_{0}}+1)},$
$q$ $=\gamma\omega(S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}})$ $\quad$=$\quad$= $\frac{1}{2(e^{\varepsilon_{0}}+1)}-\frac{2}{|\mathcal{H}|(1+e^{\varepsilon_{0}})}$ $\quad$=$\quad$= $\frac{1}{2(e^{\varepsilon_{0}}+1)}-\frac{1}{2^{D-1}(1+e^{\varepsilon_{0}})},$
$r$ $=\gamma\omega(S_{1,1})$ $\quad$=$\quad$= $\frac{1}{2(e^{\varepsilon_{0}}+1)}+\frac{2e^{\varepsilon_{0}}}{|\mathcal{H}|(1+e^{\varepsilon_{0}})}$ $\quad$=$\quad$= $\frac{1}{2(e^{\varepsilon_{0}}+1)}+\frac{e^{\varepsilon_{0}}}{2^{D-1}(1+e^{\varepsilon_{0}})}.$
@@
($D$ denotes the size of the domain.) $p$ $q$ $r$ $k$-RR [War65] $\frac{1}{e^{\varepsilon_{0}}+k-1}$ 00 $\frac{k-2}{e^{\varepsilon_{0}}+k-1}$ BLH [WBLJ17] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}-\frac{1}{2^{D-1}(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}+\frac{e^{\varepsilon_{0}}}{2^{D-1}(e^{\varepsilon_{0}}+1)}$ RAPPOR [EPK14] $\frac{1}{(e^{\varepsilon_{0}/2}+1)^{2}}$ $\frac{e^{-\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{2}}-\frac{e^{-\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{D}}$ $\frac{e^{\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{2}}+\frac{e^{\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{D}}$ OUE [WBLJ17] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{e^{-\varepsilon_{0}}}{2(e^{\varepsilon_{0}}+1)}-\frac{e^{-\varepsilon_{0}}}{2(e^{\varepsilon_{0}}+1)^{D-1}}$ $\frac{e^{\varepsilon_{0}}}{2(e^{\varepsilon_{0}}+1)}+\frac{1}{2(e^{\varepsilon_{0}}+1)^{D-1}}$ HR [ASZ19] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$
Table 3: Optimal decompositions for common local randomizers (Asymptotic)
@@
BLH ($D\gg 1$[WBLJ17] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$
RAPPOR ($D\gg 1$[EPK14] $\frac{1}{(e^{\varepsilon_{0}/2}+1)^{2}}$ $\frac{1}{e^{\varepsilon_{0}/2}(e^{\varepsilon_{0}/2}+1)^{2}}$ $\frac{e^{\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{2}}$
OUE ($D\gg 1$[WBLJ17] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2e^{\varepsilon_{0}}(e^{\varepsilon_{0}}+1)}$ $\frac{e^{\varepsilon_{0}}}{2(e^{\varepsilon_{0}}+1)}$
@@ A.2 RAPPOR
where $d_{H}(\mathrm{UE}(x),\boldsymbol{y})$ denotes the Hamming distance between two binary vectors in $\{0,1\}^{D}$ [EPK14].
@@
$S_{e^{\varepsilon_{0}},1}$ =$=\{\boldsymbol{y}\mid\boldsymbol{y}[x_{1}^{0}]=1\wedge\boldsymbol{y}[x_{1}^{1}]=0\},$
$S_{1,e^{\varepsilon_{0}}}$ =$=\{\boldsymbol{y}\mid\boldsymbol{y}[x_{1}^{0}]=0\wedge\boldsymbol{y}[x_{1}^{1}]=1\},$
$S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}}$ =$=\{\boldsymbol{y}\mid\boldsymbol{y}[x_{1}^{0}]=1\wedge\boldsymbol{y}[x_{1}^{1}]=1\}-\{\boldsymbol{y}^{*}\},$
$S_{1,1}$ =$=\{\boldsymbol{y}\mid\boldsymbol{y}[x_{1}^{0}]=0\wedge\boldsymbol{y}[x_{1}^{1}]=0\}\cup\{\boldsymbol{y}^{*}\}.$
It can be computed that
$p$ $=\gamma\omega(S_{e^{\varepsilon_{0}},1})$ $\quad$=$\quad$= $\frac{1}{(e^{\varepsilon_{0}/2}+1)^{2}},$
$q$ $=\gamma\omega(S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}})$ $\quad$=$\quad$= $\frac{e^{-\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{2}}-\frac{e^{-\varepsilon_{0}/2}}{(1+e^{\varepsilon_{0}/2})^{D}},$
$r$ $=\gamma\omega(S_{1,1})$ $\quad$=$\quad$= $\frac{e^{\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{2}}+\frac{e^{\varepsilon_{0}/2}}{(1+e^{\varepsilon_{0}/2})^{D}}.$
@@ A.3 Optimal Unary Encode (OUE)
OUE (Optimized Unary Encoding) follows the same initial step as RAPPOR by applying unary encoding to the input $x$, resulting in $\mathrm{UE}(x)\in\{0,1\}^{D}$. However, it differs in how randomized response is applied: for each bit in $\mathrm{UE}(x)$, a 2-ary randomized response mechanism satisfying $\varepsilon_{0}$-DP is applied to the bits with value 0, while the bits with value 1 are replaced with uniformly random bits [WBLJ17].
@@
$S_{e^{\varepsilon_{0}},1}$ =$=\{\boldsymbol{y}\mid\boldsymbol{y}[x_{1}^{0}]=1\wedge\boldsymbol{y}[x_{1}^{1}]=0\},$
$S_{1,e^{\varepsilon_{0}}}$ =$=\{\boldsymbol{y}\mid\boldsymbol{y}[x_{1}^{0}]=0\wedge\boldsymbol{y}[x_{1}^{1}]=1\},$
$S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}}$ =$=\{\boldsymbol{y}\mid\boldsymbol{y}[x_{1}^{0}]=1\wedge\boldsymbol{y}[x_{1}^{1}]=1\}-\{\boldsymbol{y}^{*}\},$
$S_{1,1}$ =$=\{\boldsymbol{y}\mid\boldsymbol{y}[x_{1}^{0}]=0\wedge\boldsymbol{y}[x_{1}^{1}]=0\}\cup\{\boldsymbol{y}^{*}\}.$
It can be computed that
$p$ $=\gamma\omega(S_{e^{\varepsilon_{0}},1})$ $\quad$=$\quad$= $\frac{1}{2(1+e^{\varepsilon_{0}})},$
$q$ $=\gamma\omega(S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}})$ $\quad$=$\quad$= $\frac{e^{-\varepsilon_{0}}}{2(1+e^{\varepsilon_{0}})}-\frac{e^{-\varepsilon_{0}}}{2(1+e^{\varepsilon_{0}})^{D-1}},$
$r$ $=\gamma\omega(S_{1,1})$ $\quad$=$\quad$= $\frac{e^{\varepsilon_{0}}}{2(e^{\varepsilon_{0}}+1)}+\frac{1}{2(1+e^{\varepsilon_{0}})^{D-1}}.$
@@ A.4 Hadamard Response
The Hadamard Response [ASZ19] assumes the input domain $\mathcal{X}=[D]$, where $D=2^{r}$ is a power of two. The mechanism maps each input $x\in[D]$ to a row of the normalized Hadamard matrix $H\in\{-1,1\}^{D\times D}$, and samples an output index $y\in[D]$ with probability proportional to $\exp\left(\frac{\varepsilon_{0}}{2}H_{x,y}\right)$, where $\varepsilon_{0}$ is the privacy budget. Each entry of the Hadamard matrix is defined as $H_{x,y}=(-1)^{x\odot y}$, where $x\odot y$ denotes the bitwise inner product modulo 2 between the binary representations of $x$ and $y$; that is, the sum (mod 2) of the component-wise products of their bits.
@@
By direct computation from the definition, the blanket distribution of the Hadamard Response is the uniform distribution over $[D]$. Its corresponding parameter is $\gamma=\frac{2}{1+e^{\varepsilon_{0}}}$.
@@
$S_{e^{\varepsilon_{0}},1}$ =$=\{y|x_{1}^{0}\odot y=1\wedge x_{1}^{1}\odot y=0\},$
$S_{1,e^{\varepsilon_{0}}}$ =$=\{y|x_{1}^{0}\odot y=0\wedge x_{1}^{1}\odot y=1\},$
$S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}}$ =$=\{y|x_{1}^{0}\odot y=1\wedge x_{1}^{1}\odot y=1\},$
$S_{1,1}$ =$=\{y|x_{1}^{0}\odot y=0\wedge x_{1}^{1}\odot y=0\}.$
@@
$p$ $=\gamma\omega(S_{e^{\varepsilon_{0}},1})$ $\quad$=$\quad$= $\frac{1}{2(1+e^{\varepsilon_{0}})},$
$q$ $=\gamma\omega(S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}})$ $\quad$=$\quad$= $\frac{1}{2(1+e^{\varepsilon_{0}})},$
$r$ $=\gamma\omega(S_{1,1})$ $\quad$=$\quad$= $\frac{1}{2(1+e^{\varepsilon_{0}})}.$

### v6

@@ 8 Limitation
However, when the adjacency relation is defined over the full product domain $\mathbb{X}^{1}\times\mathbb{X}^{2}\times\dots\times\mathbb{X}^{m}$, one must also consider neighboring pairs that differ in only a subset of coordinates. If all $\mathcal{R}_{i}$ are identical, we need to separately consider adjacency relations with Hamming distances $1,2,\dots,m$. For example, in the joint composition of three $\varepsilon_{0}/4$-DP $k$-$R$ mechanisms, we must evaluate the GPARV for the following input pairs:
$\big((0,0,0,0),(1,1,1,1)\big),\quad\big((0,0,0,0),(0,1,1,1)\big),\quad\big((0,0,0,0),(0,0,1,1)\big),\quad\big((0,0,0,0),(0,0,0,1)\big).$
@@ A.1 Binary local hash
Let $\mathcal{X}=[D]$. The Binary Local Hashing (BLH) mechanism is defined as $\mathsf{BLH}(x)=(h,\mathsf{RR}(h(x)))$, where $h:[D]\to[2]$ is a uniformly random function drawn from $\mathcal{H}=\{f:[D]\to[2]\}$, and $\mathsf{RR}$ denotes the 2-ary randomized response mechanism (see Example 4.1 for the definition) [3].
@@
In summary, the blanket distribution of BLH is “nearly” uniform distribution over $\mathcal{H}\times[2]$:
@@
$S_{r_{0},r_{1}}=\{y|\frac{\mathcal{R}(x_{1}^{0})(y)}{\inf_{x}\mathcal{R}(x)(y)}=r_{0},\frac{\mathcal{R}(x_{1}^{1})(y)}{\inf_{x}\mathcal{R}(x)(y)}=r_{1}\}.$
@@
$S_{e^{\varepsilon_{0}},1}$ $=\{(h,y)|h(x_{1}^{0})=y\wedge h(x_{1}^{1})\neq y\},$
$S_{1,e^{\varepsilon_{0}}}$ $=\{(h,y)|h(x_{1}^{0})\neq y\wedge h(x_{1}^{1})=y\},$
$S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}}$ $=\{(h,y)|h(x_{1}^{0})=y\wedge h(x_{1}^{1})=y\}-\{(h,y)|h=h^{*}\vee h=\bar{h}^{*}\},$
$S_{1,1}$ $=\{(h,y)|h(x_{1}^{0})\neq y\wedge h(x_{1}^{1})\neq y\}\cup\{(h,y)|h=h^{*}\vee h=\bar{h}^{*}\}.$
@@
$p$ $=\gamma\omega(S_{e^{\varepsilon_{0}},1})$ $\quad=\quad$ $\frac{1}{2(e^{\varepsilon_{0}}+1)},$
$q$ $=\gamma\omega(S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}})$ $\quad=\quad$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}-\frac{2}{|\mathcal{H}|(1+e^{\varepsilon_{0}})}$ $\quad=\quad$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}-\frac{1}{2^{D-1}(1+e^{\varepsilon_{0}})},$
$r$ $=\gamma\omega(S_{1,1})$ $\quad=\quad$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}+\frac{2e^{\varepsilon_{0}}}{|\mathcal{H}|(1+e^{\varepsilon_{0}})}$ $\quad=\quad$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}+\frac{e^{\varepsilon_{0}}}{2^{D-1}(1+e^{\varepsilon_{0}})}.$
@@
($D$ denotes the size of the domain.) $p$ $q$ $r$ $k$-$R$ [11] $\frac{1}{e^{\varepsilon_{0}}+k-1}$ 0 $\frac{k-2}{e^{\varepsilon_{0}}+k-1}$ BLH [3] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}-\frac{1}{2^{D-1}(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}+\frac{e^{\varepsilon_{0}}}{2^{D-1}(e^{\varepsilon_{0}}+1)}$ RAPPOR [6] $\frac{1}{(e^{\varepsilon_{0}/2}+1)^{2}}$ $\frac{e^{-\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{2}}-\frac{e^{-\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{D}}$ $\frac{e^{\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{2}}+\frac{e^{\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{D}}$ OUE [3] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{e^{-\varepsilon_{0}}}{2(e^{\varepsilon_{0}}+1)}-\frac{e^{-\varepsilon_{0}}}{2(1+e^{\varepsilon_{0}})^{D-1}}$ $\frac{e^{\varepsilon_{0}}}{2(e^{\varepsilon_{0}}+1)}+\frac{1}{2(e^{\varepsilon_{0}}+1)^{D-1}}$ HR [HR] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}+\frac{2(e^{\varepsilon_{0}}-1)}{D(e^{\varepsilon_{0}}+1)}$
[…]
@@
BLH  [3] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$
RAPPOR  [6] $\frac{1}{(e^{\varepsilon_{0}/2}+1)^{2}}$ $\frac{1}{e^{\varepsilon_{0}/2}(e^{\varepsilon_{0}/2}+1)^{2}}$ $\frac{e^{\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{2}}$
OUE  [3] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2e^{\varepsilon_{0}}(e^{\varepsilon_{0}}+1)}$ $\frac{e^{\varepsilon_{0}}}{2(e^{\varepsilon_{0}}+1)}$
HR  [HR] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$
@@ A.2 RAPPOR
where $d_{H}(\mathrm{UE}(x),\boldsymbol{y})$ denotes the Hamming distance between two binary vectors in $\{0,1\}^{D}$.
@@
$S_{e^{\varepsilon_{0}},1}$ $=\{\boldsymbol{y}|\boldsymbol{y}[x_{1}^{0}]=1\wedge\boldsymbol{y}[x_{1}^{1}]=0\},$
$S_{1,e^{\varepsilon_{0}}}$ $=\{\boldsymbol{y}|\boldsymbol{y}[x_{1}^{0}]=0\wedge\boldsymbol{y}[x_{1}^{1}]=1\},$
$S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}}$ $=\{\boldsymbol{y}|\boldsymbol{y}[x_{1}^{0}]=1\wedge\boldsymbol{y}[x_{1}^{1}]=1\}-\{\boldsymbol{y}^{*}\},$
$S_{1,1}$ $=\{\boldsymbol{y}|\boldsymbol{y}[x_{1}^{0}]=0\wedge\boldsymbol{y}[x_{1}^{1}]=0\}\cup\{\boldsymbol{y}^{*}\}.$
When $D>2$, it can be computed that
$p$ $=\gamma\omega(S_{e^{\varepsilon_{0}},1})$ $\quad=\quad$ $\frac{1}{(e^{\varepsilon_{0}/2}+1)^{2}},$
$q$ $=\gamma\omega(S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}})$ $\quad=\quad$ $\frac{e^{-\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{2}}-\frac{e^{-\varepsilon_{0}/2}}{(1+e^{\varepsilon_{0}/2})^{D}},$
$r$ $=\gamma\omega(S_{1,1})$ $\quad=\quad$ $\frac{e^{\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{2}}+\frac{e^{\varepsilon_{0}/2}}{(1+e^{\varepsilon_{0}/2})^{D}}.$
@@ A.3 Optimal Unary Encode (OUE)
OUE (Optimized Unary Encoding) follows the same initial step as RAPPOR by applying unary encoding to the input $x$, resulting in $\mathrm{UE}(x)\in\{0,1\}^{D}$. However, it differs in how randomized response is applied: for each bit in $\mathrm{UE}(x)$, a 2-ary randomized response mechanism satisfying $\varepsilon_{0}$-DP is applied to the bits with value 0, while the bits with value 1 are replaced with uniformly random bits [3].
@@
$S_{e^{\varepsilon_{0}},1}$ $=\{\boldsymbol{y}|\boldsymbol{y}[x_{1}^{0}]=1\wedge\boldsymbol{y}[x_{1}^{1}]=0\},$
$S_{1,e^{\varepsilon_{0}}}$ $=\{\boldsymbol{y}|\boldsymbol{y}[x_{1}^{0}]=0\wedge\boldsymbol{y}[x_{1}^{1}]=1\},$
$S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}}$ $=\{\boldsymbol{y}|\boldsymbol{y}[x_{1}^{0}]=1\wedge\boldsymbol{y}[x_{1}^{1}]=1\}-\{\boldsymbol{y}^{*}\},$
$S_{1,1}$ $=\{\boldsymbol{y}|\boldsymbol{y}[x_{1}^{0}]=0\wedge\boldsymbol{y}[x_{1}^{1}]=0\}\cup\{\boldsymbol{y}^{*}\}.$
When $D>2$, it can be computed that
$p$ $=\gamma\omega(S_{e^{\varepsilon_{0}},1})$ $\quad=\quad$ $\frac{1}{2(1+e^{\varepsilon_{0}})},$
$q$ $=\gamma\omega(S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}})$ $\quad=\quad$ $\frac{e^{-\varepsilon_{0}}}{2(1+e^{\varepsilon_{0}})}-\frac{e^{-\varepsilon_{0}}}{2(1+e^{\varepsilon_{0}})^{D-1}},$
$r$ $=\gamma\omega(S_{1,1})$ $\quad=\quad$ $\frac{e^{\varepsilon_{0}}}{2(e^{\varepsilon_{0}}+1)}+\frac{1}{2(1+e^{\varepsilon_{0}})^{D-1}}.$
@@ A.4 Hadamard Response
The Hadamard Response [HR] assumes an input domain $\mathbb{X}=[D]$, where $D=2^{r}$ is a power of two [^3] Because of properties of the Hadamard matrix, 0 is treated as a special symbol and is excluded from the input domain.. The mechanism maps each input $x\in[D]$ to a row of the normalized Hadamard matrix $H\in\{-1,1\}^{D\times D}$, and samples an output index $y\in[D]$ with probability proportional to $\exp\left(\frac{\varepsilon_{0}}{2}H_{x,y}\right)$, where $\varepsilon_{0}$ is the privacy budget. Each entry of the Hadamard matrix is defined as $H_{x,y}=(-1)^{x\odot y}$, where $x\odot y$ denotes the bitwise inner product modulo 2 between the binary representations of $x$ and $y$; that is, the sum (mod 2) of the component-wise products of their bits.
@@
Similarly, HR also exhibits an extreme case:
$\inf_{x\in\mathbb{X}}\Pr[\mathcal{R}(x)=0]$ $=\frac{2e^{\varepsilon_{0}}}{(1+e^{\varepsilon_{0}})D},$
$\forall y\neq 0,\inf_{x\in\mathbb{X}}\Pr[\mathcal{R}(x)=y]$ $=\frac{2}{(1+e^{\varepsilon_{0}})D}.$
By direct computation from the definition, its corresponding parameter is $\gamma=\frac{2}{1+e^{\varepsilon_{0}}}+\frac{2(e^{\varepsilon_{0}}-1)}{(1+e^{\varepsilon_{0}})D}$.
@@
$S_{e^{\varepsilon_{0}},1}$ $=\{y|x_{1}^{0}\odot y=1\wedge x_{1}^{1}\odot y=0\},$
$S_{1,e^{\varepsilon_{0}}}$ $=\{y|x_{1}^{0}\odot y=0\wedge x_{1}^{1}\odot y=1\},$
$S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}}$ $=\{y|x_{1}^{0}\odot y=1\wedge x_{1}^{1}\odot y=1\},$
$S_{1,1}$ $=\{y|x_{1}^{0}\odot y=0\wedge x_{1}^{1}\odot y=0\}.$
@@
$p$ $=\gamma\omega(S_{e^{\varepsilon_{0}},1})$ $\quad=\quad$ $\frac{1}{2(1+e^{\varepsilon_{0}})},$
$q$ $=\gamma\omega(S_{e^{\varepsilon_{0}},e^{\varepsilon_{0}}})$ $\quad=\quad$ $\frac{1}{2(1+e^{\varepsilon_{0}})},$
$r$ $=\gamma\omega(S_{1,1})$ $\quad=\quad$ $\frac{1}{2(1+e^{\varepsilon_{0}})}+\frac{2(e^{\varepsilon_{0}}-1)}{(1+e^{\varepsilon_{0}})D}.$